OpenStudy (anonymous):

mike as 8 more goldfish than alan, alan has 4 fewer goldfish than susie. mike has 12 fish. how many altogether?

OpenStudy (anonymous):

12-8=4 4+4=8 12+4+8=24 goldfish

OpenStudy (anonymous):

right? medal?

OpenStudy (anonymous):

thank you!